How do I work out the WACC

The company’s preference shares are currently trading at $0.70 each. The company’s ordinary shares are currently trading at $1.55 each.

• The risk-free rate of return is 2.36 % p.a., and the return on the market is 8.50 % p.a.

• Debentures have a coupon interest rate of 9% p.a. and could be re-issued at the present time at an interest rate of 7% p.a. The debentures will be redeemed at their face value in three years’ time. Face value is as per the balance sheet.

• The mortgage loan is repayable in seven years’ time and the current interest rate is 4.85% p.a. The mortgage was initially negotiated at 8.15% p.a.

• Term loans have a current interest rate of 5% p.a., but were negotiated at an interest rate of 6% p.a. They are repayable in full in three years’ time. The term loans consists of regular semi-annually interest payments with the principal repaid at maturity

• Unsecured notes will mature in six months and will not be replaced. They have a current interest rate of 2.74% p.a.

• The current interest rate on the bank overdraft is 5.0% p.a.

• Interest on all debt securities is paid twice-yearly and the corporate tax-rate is 32 percent

Balance Sheet for Kidman Resources

Issued Capital $ million

308,000,000 Ordinary shares of $2.50 full paid 770.00 2,500,000 3.5% Preference Shares of $1.00 fully paid 2.50

Current Liabilities and Provisions

Bank Overdraft 3.00

Trade Creditors 5.00

Unsecured Notes 6.00

Non-Current Liabilities

Debentures 14.00

Term Loans 7.00

Mortgage 15.00
Answer #1
Cost of Equity = Risk free rate + Beta * (Market return - Risk free return)
=2.36+(1*(8.50-2.36)
8.5
Assuming beta is to be 1 as beta is not provided
A B A*B
Source Amount % of total Capital Pre tax cost Cost*(1-tax rate) Post tax cost Cost of Capital
Bank Overdraft 3.00 0.37% 5.00% 5.00%*(1-0.32) = 3.40% 0.37%*3.40% = 0.01%
Unsecured Notes 6.00 0.73% 1.37% 2.74%*(1-0.32) = 0.93% 0.73%*0.93% = 0.01%
Debentures 14.00 1.71% 7.00% 7.00%*(1-0.32) = 4.76% 1.71%*4.76% = 0.08%
Term Loans 7.00 0.86% 5.00% 5.00%*(1-0.32) = 3.40% 0.86%*3.40% = 0.03%
Mortgage 15.00 1.83% 4.85% 4.85%*(1-0.32) = 3.30% 1.83%*3.30% = 0.06%
Preference shares 2.50 0.31% 3.50% 3.50% 0.31%*3.50% = 0.01%
Ordinary shares 770.00 94.19% 8.50% 8.50% 94.19%*8.50% = 8.01%
Total Capital 817.50 WACC 8.21%